# Large Vocabulary Size Improves Large Language Models

### v1

@@ 3.1 Settings
For example, with a fixed number of training tokens, the 500k vocabulary model trains for around 1.5 epochs in English and 2 epochs in Japanese, while the 5k vocabulary model trains for only 1 epoch. The larger vocabulary size has an advantage of seeing more data in this configuration. In contrast, with a fixed number of training epochs, the 5k vocabulary model consumes much more computational resources than the larger vocabulary models. Especially in Japanese, where the 5k vocabulary model contains about twice as much tokens as the 500k vocabulary model in 1 epoch. Because the performance of LLMs is correlated with the computational costs during training Kaplan et al. (2020), this configuration might favor smaller vocabulary sizes. Thus, we prepare two training configurations: 1T tokens and 1 epoch [^1] In addition to the training data size, we have to discuss the number of parameters for a fair comparison because the model with the small vocabulary size contains less parameters for the embedding and output layers. However, as described in Appendix B, the model with the small vocabulary size does not improve the performance when we increase the number of parameters related to the vocabulary. Thus, we focus only on varying the training data size in our main experiments..
For hyper-parameters of the language model, we use the GPT-3 Large setting described in Brown et al. (2020). We set the number of layers $24$ and the hidden dimension size $1536$. In this setting, the number of parameters for internal layers is 680M [^2] We describe details of hyper-parameters in Appendix A.. We use Megatron-LM Shoeybi et al. (2020) [^3] https://github.com/NVIDIA/Megatron-LM as our codebase to train large language models. To stabilize the training, we use the scaled embed technique as described by Takase et al. (2024).
@@ 3.2 Results
Table 2: The performance on English commonsense reasoning tasks when we use 1T tokens for training.
@@
Table 3: The performance on Japanese commonsense reasoning tasks when we use 1T tokens for training.
#Vocab PIQA OBQA HellaSwag WinoGrande ARC-e ARC-c Avg.
5k 70.1 32.4 50.9 55.2 50.2 28.5 47.9 (±$\pm 0.0$)
10k 70.1 32.0 50.7 53.6 50.3 26.8 47.2 ($-0.7$)
50k 70.6 33.6 52.1 53.8 52.3 27.3 48.3 ($+0.4$)
100k 71.7 33.8 53.4 54.7 52.7 27.6 49.0 ($+1.1$)
500k 70.4 34.2 54.3 55.1 54.0 28.2 49.4 ($+1.5$)
Table 4: The performance on English commonsense reasoning tasks when we train 1 epoch for each training data.
#Vocab JSQuAD JCQA XWinograd JAQKET Avg.
@@
Table 5: The performance on Japanese commonsense reasoning tasks when we train 1 epoch for each training data.
Tables 2 and 3 present the performance of the models trained with 1T tokens, and Tables 4 and 5 show the performance of the models trained for 1 epoch. For each configuration, we show the average score of each task, and the improvement of the average score from the 5k vocabulary model.
@@ 4 Experiments on Continual Training
4.1 Increasing Vocabulary Size in Continual Training
@@
Let $V_{orig}$ and $V_{new}$ be the vocabulary sizes of the pre-trained model and a newly constructed vocabulary respectively, and let $d$ be the dimension size of each layer. To exploit knowledge learned in the pre-trained embedding matrix, we construct a new embedding matrix $E_{new}\in\mathbb{R}^{V_{new}\times d}$ from the original embedding matrix $E_{orig}\in\mathbb{R}^{V_{orig}\times d}$ as follows:
$E_{new}=\frac{WE_{orig}}{\sqrt{V_{orig}}},$ (1)
where $W\in\mathbb{R}^{V_{new}\times V_{orig}}$ is the random matrix whose elements are sampled from the standard normal distribution independently. To maintain the standard deviation of $E_{orig}$ in $E_{new}$, we scale the matrix multiplication by $\frac{1}{\sqrt{V_{orig}}}$ [^7] We assume that $E_{orig}$ contains independent random variables with mean 00 and variance $\mathrm{var}(E_{orig})$. Then, the variance of the matrix multiplication $WE_{orig}$ has mean 00 and variance $\mathrm{var}(E_{orig})\times V_{orig}$..
In the naive way, we swap $E_{new}$ with $E_{orig}$. However, Equation 1 randomizes embeddings even if $V_{new}$ contains the corresponding subwords which may possess useful knowledge transferable to the new model. Therefore, we insert the pre-trained embedding in $E_{orig}$ into $E_{new}$ if the corresponding subword is included in both $V_{orig}$ and $V_{new}$. For the output layer, we construct a new weight matrix with the same manner.
@@ 4.2 Results
We train the Llama2 7B parameter model Touvron et al. (2023) with 100B tokens on our Japanese training data. We use the Japanese vocabulary whose size is 100k. Table 6 shows results on Japanese commonsense reasoning tasks. In this table, ‘Swap’ uses new parameters related to the vocabulary without inserting the corresponding pre-trained parameters. Moreover, we train a language model from scratch to compare the effectiveness of the continual training.
Table 6 shows that ‘Swap’ and ‘Swap&Insert’ outperform the model using the original Llama2 vocabulary even though these settings randomize parameters related to the vocabulary. This result indicates that it is better to prepare an appropriate vocabulary even in the continual training situation. Moreover, the insertion strategy achieves further improvement. Therefore, we also have to use the pre-trained embeddings if the corresponding subwords are included in $V_{orig}$ and $V_{new}$.
@@ 5 Related Work
Before the paradigm of subwords, researchers sometimes needed to handle the large vocabulary size such as more than 100k to decrease the number of unknown words. For example, the vocabulary sizes of One Billion Word Benchmark and WikiText-103 are about 800k and 300k respectively Chelba et al. (2013); Merity et al. (2017). Some previous studies reported that character-level information was useful for neural language models with the large vocabulary size Jozefowicz et al. (2016); Takase et al. (2019).
Since the use of subword units is proposed Sennrich et al. (2016); Kudo (2018), the vocabulary sizes 30k-60k are widely used as the magic numbers (Libovický et al., 2022). As examples, the BERT and GPT papers use 30k and 40k for their vocabulary sizes respectively without any justification Vaswani et al. (2017); Devlin et al. (2019); Radford et al. (2018a). Although Kiyono et al. (2020) reported that the encoder-decoder model with the vocabulary size 32k outperforms ones with smaller vocabulary sizes in the machine translation task, they did not explore the performance with larger vocabulary sizes than 32k.
@@ 6 Conclusion
In this study, we conducted experiments on two languages: English and Japanese. We believe that our findings can be applied to other languages because we do not depend on linguistic features in the subword vocabulary construction. However, we also agree that it is better to conduct exhaustive experiments on various languages to confirm the generality of our findings. Moreover, the parameter sizes of internal layers are 680M in training from scratch, and 7B in the continual training. We consider that the discussions on subword vocabulary size are orthogonal to the parameter size of internal layers, but we would conduct additional experiments with more than 10B parameters if we had a large amount of computational resources.
@@ Appendix A Hyper-parameters
Appendix B Comparison on Parameter Size
@@
The smaller vocabulary size lessens the parameter sizes related to the vocabulary in comparison with the larger vocabulary size. Thus, the smaller vocabulary size might have the disadvantage in the number of parameters. To confirm this point, we increase the parameters related to the vocabulary for the 5k setting. Concretely, we expand the dimension of the embedding and output layers, and then modify the dimension size by the linear transformation such as the matrix factorization technique Lan et al. (2020) [^8] In contrast, we can reduce the number of parameters for the larger vocabulary size with the matrix factorization technique or more sophisticated way Takase and Kobayashi (2020), but we regard the 5k as the baseline in this experiment.. Let $V$ be the vocabulary size, $d_{e}$ be the dimension size of the embedding and output layers, and $d$ be the hidden dimension size. We prepare the expanded embedding layer $E\in\mathbb{R}^{V\times d_{e}}$ and the trainable weight matrix $W\in\mathbb{R}^{d_{e}\times d}$. We convert the dimension size of $E$ with the matrix multiplication $EW$. For the output layer, we convert the dimension with the same manner. We adjust $d_{e}$ for a fair comparison with the 100k setting.
@@
Table 8 shows the performance on Japanese commonsense reasoning tasks. This table indicates that the 5k with the expansion does not improve the average score although it increases the number of parameters. This result suggests that the increase of the parameter size related to the vocabulary has no positive influence on the performance. In contrast, the 100k achieves much better average score in the similar parameter size.

### v2

@@ 3.1 Settings
For example, with a fixed number of training tokens, the 500k vocabulary model trains for around 1.5 epochs in English and 2 epochs in Japanese, while the 5k vocabulary model trains for only 1 epoch. The larger vocabulary size has an advantage of seeing more data in this configuration. In contrast, with a fixed number of training epochs, the 5k vocabulary model consumes much more computational resources than the larger vocabulary models. Especially in Japanese, where the 5k vocabulary model contains about twice as much tokens as the 500k vocabulary model in 1 epoch. Because the performance of LLMs is correlated with the computational costs during training Kaplan et al. (2020), this configuration might favor smaller vocabulary sizes. Thus, we prepare two training configurations: 1T tokens and 1 epoch [^1] In addition to the training data size, we have to discuss the number of parameters for a fair comparison because the model with the small vocabulary size contains less parameters for the embedding and output layers. However, as described in Appendix D, the model with the small vocabulary size does not improve the performance when we increase the number of parameters related to the vocabulary. Thus, we focus only on varying the training data size in our main experiments..
For hyper-parameters of the language model, we use the GPT-3 Large setting described in Brown et al. (2020). We set the number of layers $24$ and the hidden dimension size $1536$. In this setting, the number of parameters for internal layers is 680M. We use Megatron-LM Shoeybi et al. (2020) [^2] https://github.com/NVIDIA/Megatron-LM as our codebase to train large language models. To stabilize the training, we use the scaled embed technique Takase et al. (2024).
@@ 3.2 Results
1 Epoch
5k 70.1 32.4 50.9 55.2 50.2 28.5 47.9 (±$\pm 0.0$)
10k 71.1 33.6 50.6 55.7 49.0 27.1 47.9 (±$\pm 0.0$)
50k 70.6 33.6 52.1 53.8 52.3 27.3 48.3 ($+0.4$)
100k 71.7 33.8 53.4 54.7 52.7 27.6 49.0 ($+1.1$)
500k 70.4 34.2 54.3 55.1 54.0 28.2 49.4 ($+1.5$)
Table 2: The performance on English commonsense reasoning tasks in training 1T tokens and 1 epoch.
@@ 1 Epoch
1 Epoch
@@
Table 3: The performance on Japanese commonsense reasoning tasks in training 1T tokens and 1 epoch.
Tables 2 and 3 present the performance of the models trained with 1T tokens and 1 epoch. For each configuration, we show the average score of each task, and the improvement of the average score from the 5k vocabulary model.
@@ 4 Experiments on Continual Training
4.1 Increasing Vocabulary Size
@@
Let $V_{orig}$ and $V_{new}$ be the vocabulary set of the pre-trained model and a newly constructed vocabulary set respectively, and let $d$ be the dimension size of each layer. To exploit knowledge learned in the pre-trained embedding matrix, we construct a new embedding matrix $E_{new}\in\mathbb{R}^{|V_{new}|\times d}$ from the original embedding matrix $E_{orig}\in\mathbb{R}^{|V_{orig}|\times d}$ with the way inspired by the randomized algorithm Halko et al. (2011):
$E_{new}=\frac{WE_{orig}}{\sqrt{|V_{orig}|}},$ (1)
where $W\in\mathbb{R}^{|V_{new}|\times|V_{orig}|}$ is the random matrix whose elements are sampled from the standard normal distribution independently. To maintain the standard deviation of $E_{orig}$ in $E_{new}$, we scale the matrix multiplication by $\frac{1}{\sqrt{|V_{orig}|}}$ [^6] We assume that $E_{orig}$ contains independent random variables with mean 00 and variance $\mathrm{var}(E_{orig})$. Then, the variance of the matrix multiplication $WE_{orig}$ has mean 00 and variance $\mathrm{var}(E_{orig})\times|V_{orig}|$..
In the naive way, we swap $E_{new}$ with $E_{orig}$. However, Equation 1 randomizes embeddings even if $V_{new}$ contains the corresponding subwords which may possess useful knowledge transferable to the new model. Therefore, we insert the pre-trained embedding in $E_{orig}$ into $E_{new}$ if the corresponding subword is included in both $V_{orig}$ and $V_{new}$ [^7] See Appendix B for more details.. For the output layer, we construct a new weight matrix with the same manner.
@@ 4.2 Results
We train the Llama2 7B parameter model Touvron et al. (2023) with 100B tokens on our Japanese training data. We use the Japanese vocabulary whose size is 100k. Table 4 shows results on Japanese commonsense reasoning tasks. In this table, ‘Swap’ uses new parameters related to the vocabulary without inserting the corresponding pre-trained parameters. We train a language model from scratch to compare the effectiveness of the continual training. Moreover, we compare the embedding initialization method by Fujii et al. (2024) because their study is the same situation: continual training of Llama2 on Japanese data.
Table 4 shows that ‘Swap’ and ‘Swap&Insert’ outperform the model using the original Llama2 vocabulary even though these settings randomize parameters related to the vocabulary. This result indicates that it is better to prepare an appropriate vocabulary even in the continual training situation. Moreover, the insertion strategy achieves further improvement. The ‘Swap&Insert’ outperforms the method of Fujii et al. (2024), which initializes an embedding of the new subword with the average of the pre-trained embeddings [^8] For the existing subwords, their method uses the pre-trained embeddings. Thus, their method is regarded as using the ‘Insert’ strategy., and thus, the ‘Swap&Insert’ is simple but effective.
@@ 5 Related Work
Before the paradigm of subword units and LLMs, researchers sometimes needed to handle the large vocabulary size such as more than 100k to decrease the number of unknown words. For example, the vocabulary sizes of One Billion Word Benchmark and WikiText-103 are about 800k and 300k respectively Chelba et al. (2013); Merity et al. (2017). Some previous studies reported that character-level information was useful for neural language models with the large vocabulary size Jozefowicz et al. (2016); Takase et al. (2019). In this paradigm, Chen et al. (2019) explored the impact of the vocabulary size.
Since the use of subword units is proposed Sennrich et al. (2016); Kudo (2018), the vocabulary sizes 30k-60k are widely used as the magic numbers (Libovický et al., 2022). As examples, the BERT and GPT papers use 30k and 40k for their vocabulary sizes respectively without any justification Vaswani et al. (2017); Devlin et al. (2019); Radford et al. (2018a). Kiyono et al. (2020) investigated the relation between the performance and the vocabulary size but the maximum vocabulary size of their investigation is too small, i.e., 32k.
@@ 6 Conclusion
In this study, we conducted experiments on two languages: English and Japanese. We believe that our findings can be applied to other languages because we do not depend on linguistic features in the subword vocabulary construction. However, we also agree that it is better to conduct exhaustive experiments on various languages to confirm the generality of our findings.
In this study, we used 500k as the maximum vocabulary size. Because it is impractical to construct a much larger vocabulary than 500k, we could not investigate the improvement by the tremendously large vocabulary size such as one million and the upper bound of the performance. The computational time of the vocabulary construction depends on the corpus size and the desired vocabulary size. We roughly estimate that the vocabulary whose size is larger than one million requires at least over a month in its construction in our environment.
Furthermore, the parameter sizes of internal layers are 680M in training from scratch, and 7B in the continual training. We consider that the discussions on subword vocabulary size are orthogonal to the parameter size of internal layers, but we would conduct additional experiments with more than 10B parameters if we had a large amount of computational resources.
@@ Appendix A Hyper-parameters
Appendix B Formula of ‘Insert’ in Section 4.1
We formulate the procedure of ‘Insert’ in Section 4.1. Let $e_{i}^{orig}$ and $e_{i}^{new}$ be the $i$-th row vectors of $E_{orig}$ and $E_{new}$, and let $w_{i}^{orig}$ and $w_{i}^{new}$ be the corresponding subwords to $e_{i}^{orig}$ and $e_{i}^{new}$. The ‘Insert’ function, $\mathrm{Insert(\cdot)}$, replaces $e_{i}^{new}$ with $e_{i}^{orig}$ when the corresponding subword is included in the original vocabulary $V_{orig}$ as follows:
$\mathrm{Insert}(e_{i}^{new})=\begin{cases}e_{j}^{orig}&\text{if $w_{i}^{new}\in V_{orig}\land w_{j}^{orig}=w_{i}^{new}$}\\ e_{i}^{new}&\text{otherwise}\end{cases}$ (2)
Therefore, the matrix contains both the randomized embeddings and the original pre-trained embeddings after the ‘Insert’ procedure. As shown in Section 4.2, this procedure leads to further improvement.
Appendix C Comparison on Vocabulary Expansion in Continual Training
Initialization Vocab type JSQuAD JCQA XWinograd JAQKET Avg.
Fujii et al. (2024) Expansion 78.8 63.5 63.6 50.1 64.0
Swap&Insert Expansion 80.7 60.6 67.4 55.9 66.2
Fujii et al. (2024) Appropriate 81.6 77.6 69.1 61.1 72.4
Swap&Insert Appropriate 81.9 80.2 69.2 61.2 73.1
Table 6: The performance on Japanese commonsense reasoning tasks in the continual training from Llama2 when we construct the 100k vocabulary with the vocabulary expansion approach and construct the appropriate 100k vocabulary to the Japanese training data.
Type Number
UTF-8 byte pieces 256
Alphabet & number (e.g., a, the, 1) 5349
Symbol (e.g., +, =, ##) 209
Others such as Japanese characters 1083
Total 6897
Table 7: The type and number of shared subword units between the original Llama2 vocabulary and appropriate vocabulary, whose size is 100k, to the Japanese data in the continual training.
In Section 4, we conduct the continual training experiment on the scenario where we construct an appropriate vocabulary to the target language. In this scenario, most subword units in the original vocabulary might be removed. In contrast, the vocabulary expansion approach maintains the whole original vocabulary because it only adds new subword units to the original vocabulary Fujii et al. (2024). We investigate which approach is empirically better in this section.
We construct 100k vocabulary with the vocabulary expansion approach, and compare it with the appropriate vocabulary used in Section 4. We apply two strategies to initialize the embedding matrix: Fujii et al. (2024) and our ‘Swap&Insert’. Table 6 shows results of the continual training from Llama2. This table indicates that using appropriate vocabulary outperforms the vocabulary expansion approach. The appropriate vocabulary contains more subword units of the target language. We consider that this property improves the performance.
Table 7 shows the shared subword units between the original Llama2 vocabulary and the appropriate vocabulary. This table indicates that the number of shared subword units is only about 7000, which is about one-fifth of the original vocabulary. Moreover, this table suggests that the original vocabulary contains few Japanese subword units because the number of the shared Japanese characters is about 1000. Therefore, it is better to construct an entirely new vocabulary that is appropriate to the target language.
For the embedding initialization methods, Table 6 shows that our ‘Swap&Insert’ achieves better averaged score than the method of Fujii et al. (2024) in the same as the results in Section 4. Thus, our approach is also more suitable in the vocabulary expansion situation.
Appendix D Comparison on Parameter Size
@@
The smaller vocabulary size lessens the parameter sizes related to the vocabulary in comparison with the larger vocabulary size. Thus, the smaller vocabulary size might have the disadvantage in the number of parameters. To confirm this point, we increase the parameters related to the vocabulary for the 5k setting. Concretely, we expand the dimension of the embedding and output layers, and then modify the dimension size by the linear transformation such as the matrix factorization technique Lan et al. (2020) [^9] In contrast, we can reduce the number of parameters for the larger vocabulary size with the matrix factorization technique or more sophisticated way Takase and Kobayashi (2020), but we regard the 5k as the baseline in this experiment.. Let $|V|$ be the vocabulary size, $d_{e}$ be the dimension size of the embedding and output layers, and $d$ be the hidden dimension size. We prepare the expanded embedding layer $E\in\mathbb{R}^{|V|\times d_{e}}$ and the trainable weight matrix $W\in\mathbb{R}^{d_{e}\times d}$. We convert the dimension size of $E$ with the matrix multiplication $EW$. For the output layer, we convert the dimension with the same manner. We adjust $d_{e}=30720$ for a fair comparison with the 100k setting in terms of the number of parameters. For other hyper-parameters, we use the values shown in Table 5.
@@
Table 8 shows the performance on Japanese commonsense reasoning tasks. This table indicates that the 5k with the expansion does not improve the average score although it increases the number of parameters. This result suggests that the increase of the parameter size related to the vocabulary has no positive influence on the performance. In contrast, the 100k achieves much better average score in the similar parameter size. Therefore, the improvement by the increase of the vocabulary size is orthogonal to the increase of the parameter size.
@@
Appendix E Experiments on Each Training Data Size
In addition to the 1T tokens in Section 3, we investigate the performance in other training data sizes: 10B, 50B, 100B, 200B, and 500B tokens. Tables 9 and 10 show the results of English and Japanese models when we use each training data size. These tables show that larger vocabulary sizes lead to better performance for both English and Japanese in all training data sizes in the same as the results in Section 3. These tables indicate that our findings are independent from the amount of training data.
The difference of the performance among vocabulary sizes is smaller in the 10B tokens than ones in other training data sizes. Thus, the small training data size decreases the advantage of the large vocabulary sizes. These results explain the relation between our findings and the previous study Ali et al. (2024). Ali et al. (2024) concluded that the small vocabulary size such as 30k is sufficient for English monolingual LLMs. We consider that they led the contrary conclusion to our findings because their training data, which is about 50B tokens, is much smaller than ours.
#Vocab PIQA OBQA HellaSwag WinoGrande ARC-e ARC-c Avg.
10B tokens
5k 58.4 25.4 29.3 51.9 34.3 22.3 36.9 (±$\pm 0.0$)
10k 59.1 27.8 29.6 53.2 35.0 21.6 37.7 ($+0.8$)
50k 62.1 26.2 29.5 49.9 38.7 21.9 38.0 ($+1.1$)
100k 62.2 27.8 29.7 49.6 39.0 22.7 38.5 ($+1.6$)
500k 62.1 27.6 30.1 51.3 38.7 22.9 38.8 ($+1.9$)
50B tokens
5k 66.7 28.0 39.0 52.3 41.9 23.8 41.9 (±$\pm 0.0$)
10k 66.3 30.4 39.5 51.0 42.6 25.3 42.5 ($+0.6$)
50k 68.1 29.4 40.9 50.9 46.9 25.5 43.6 ($+1.7$)
100k 68.1 31.6 42.0 51.3 46.9 25.5 44.2 ($+2.3$)
500k 68.8 32.2 43.1 52.0 47.9 25.7 44.9 ($+3.0$)
100B tokens
5k 67.2 30.8 42.7 52.2 44.1 26.7 44.0 (±$\pm 0.0$)
10k 68.9 31.6 42.7 51.6 45.1 25.7 44.3 ($+0.3$)
50k 68.9 30.8 45.1 52.6 49.1 26.2 45.5 ($+1.5$)
100k 70.2 31.6 46.1 52.9 49.1 25.8 45.9 ($+1.9$)
500k 70.4 31.6 47.0 53.0 50.0 28.2 46.7 ($+2.7$)
200B tokens
5k 68.8 32.8 45.3 53.4 46.0 25.3 45.2 (±$\pm 0.0$)
10k 69.0 31.6 46.2 53.3 45.7 26.5 45.4 ($+0.2$)
50k 70.5 31.0 47.9 53.8 50.0 26.1 46.6 ($+1.4$)
100k 70.5 33.6 49.2 54.6 50.6 26.2 47.4 ($+2.2$)
500k 70.7 33.4 50.2 54.3 51.8 29.6 48.3 ($+3.1$)
500B tokens
5k 69.7 32.6 49.6 52.9 47.7 26.4 46.5 (±$\pm 0.0$)
10k 70.8 34.2 49.7 54.5 49.0 26.2 47.4 ($+0.9$)
50k 70.2 32.2 52.0 54.4 51.6 27.2 47.9 ($+1.4$)
100k 70.1 33.4 52.7 55.3 52.8 27.8 48.7 ($+2.2$)
500k 71.1 31.8 53.6 56.5 53.9 28.8 49.3 ($+2.8$)
Table 9: The performance on English commonsense reasoning tasks when we use 100B, 200B, and 500B tokens for training.
#Vocab JSQuAD JCQA XWinograd JAQKET Avg.
10B tokens
5k 1.6 37.1 51.0 0.9 22.7 (±$\pm 0.0$)
10k 1.4 44.2 53.6 0.5 24.9 ($+2.2$)
50k 2.7 47.9 51.0 1.6 25.8 ($+3.1$)
100k 5.3 48.9 51.7 3.3 27.3 ($+4.6$)
500k 10.1 50.8 52.5 4.1 29.4 ($+6.7$)
50B tokens
5k 36.3 49.4 53.7 3.3 35.7 (±$\pm 0.0$)
10k 42.6 59.1 56.0 7.7 41.4 ($+5.7$)
50k 42.8 56.8 55.7 12.2 41.9 ($+6.2$)
100k 40.9 56.8 56.9 17.5 43.0 ($+7.3$)
500k 48.9 57.2 54.8 17.9 44.7 ($+9.0$)
100B tokens
5k 45.0 55.0 56.9 5.2 40.5 (±$\pm 0.0$)
10k 49.8 60.9 56.7 12.3 44.9 ($+4.4$)
50k 51.4 56.0 56.8 18.4 45.7 ($+5.2$)
100k 49.1 58.7 58.9 20.7 46.9 ($+6.4$)
500k 56.3 60.1 55.7 26.3 49.6 ($+9.1$)
200B tokens
5k 50.5 58.5 56.7 7.5 43.3 (±$\pm 0.0$)
10k 53.8 61.1 57.7 16.9 47.4 ($+4.1$)
50k 55.8 54.0 58.7 21.4 47.5 ($+4.2$)
100k 54.7 64.2 58.0 27.2 51.0 ($+7.7$)
500k 60.6 61.4 56.4 32.1 52.6 ($+9.3$)
500B tokens
5k 56.4 62.2 58.4 10.4 46.9 (±$\pm 0.0$)
10k 59.6 62.8 58.8 20.2 50.4 ($+3.5$)
50k 59.3 64.7 59.0 26.5 52.4 ($+5.5$)
100k 60.1 64.7 59.3 31.8 54.0 ($+7.1$)
500k 62.6 62.9 58.8 36.4 55.2 ($+8.3$)
Table 10: The performance on Japanese commonsense reasoning tasks when we use 100B, 200B, and 500B tokens for training.